# Nonseparability of multipartite systems in dilaton black hole

### v1

@@ I Introduction
From Einstein’s perspective, the gravitational collapse of sufficiently massive stars leads to the formation of black holes—fascinating and profound entities that challenge and enrich our understanding of the universe. In recent years, both indirect and direct astronomical observations have provided strong evidence confirming the existence of these mysterious cosmic objects Q3 ; Q4 ; Q5 ; Q6 ; Q7 ; Q8 ; Q9 ; Q10 ; Q11 . A prominent example of a black hole within this framework is the Garfinkle-Horowitz-Strominger (GHS) dilaton black hole, a solution that arises from the intriguing predictions of string theory rfv01 ; rfv02 ; rfv1 ; rfv2 ; rfv3 . The presence of a dilaton field can significantly alter the properties of black hole geometries, making the study of dilaton black holes crucial for both gravitational theory and quantum mechanics. Due to their immense distances from us and their extraordinary properties, black holes have long remained enveloped in mystery, becoming a central focus of cutting-edge research. One of the most intriguing challenges in this field is the information loss paradox of the black hole Q12 ; rfv4 ; rfv5 ; rfv6 . As we all know, Hawking radiation is believed to be the root cause of the information loss paradox, as it suggests that information falling into a black hole may be irretrievably lost, seemingly violating the principles of quantum mechanics. Quantum properties, particularly in the context of Hawking radiation, may offer a crucial pathway toward resolving this paradox. Moreover, in realistic environments, quantum systems are inevitably influenced by gravitational effects, making it essential to consider the impact of Hawking radiation on quantum resources.
@@
Quantum information in gravitational background R1 ; R2 ; R3 ; R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; BJ1 ; BJ2 ; BJ3 ; R10 ; R11 ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R18 ; RM1 ; RM2 ; RM3 ; RM4 ; RM5 ; RM6 ; RM7 ; RM8 ; RM9 ; RM10 ; RM11 ; RM12 ; QQRM12 ; WDD1 ; WDD2 ; WDD3 ; BJ4 ; WDD4 ; WDD5 ; B31 ; QQQQR8 ; qert1 , an interdisciplinary field that bridges quantum information theory, quantum field theory, and general relativity. Theoretically, the influence of the gravitational effect on bipartite steering, entanglement, discord, coherence, quantum Fisher information, and entropic uncertainty relations has been investigated extensively in the background of a dilaton black hole A1 ; A2 ; A3 ; A4 ; A5 ; A6 ; A7 ; A8 ; A9 ; A10 ; A11 ; A12 ; A13 ; A14 ; A15 . As the complexity of quantum information tasks continues to increase, bipartite and tripartite inseparable states can no longer meet the growing demands, necessitating the use of N-partite inseparable states to handle these advanced tasks. For example, Deng et al. have reported new Gaussian boson sampling experiments with pseudo-photon-number-resolving detection, which register up to 255 photon-click events Q13 , and Google quantum AI and collaborators have demonstrated surface code memory below threshold in new Willow architecture by using 105 nonseparable qubits Q14 . It is necessary to understand the N-partite nonseparable quantum systems in gravitational background, because the influence of gravitational effects on realistic quantum systems is a non-negligible factor. However, measuring N-partite entanglement and obtaining its analytical expression in the dilaton black hole, particularly for complex W state, is a challenging task. Thus, one of our primary motivations is to use the Abe-Rajagopal (AR) $q$-conditional entropy to quantify the nonseparability of N-partite systems under the dilaton black hole background. Notably, this approach places stricter constraints on separability compared to traditional methods, such as the von Neumann conditional entropy, offering more nuanced insights into quantum correlations under the extreme gravitational background A17 ; A18 ; A19 .
@@
Previous studies have indicated that fermionic coherence is generally weaker than bosonic coherence in the dilaton black hole background A11 ; A16 . However, the precise differences between the nonseparability of bosonic and fermionic states in dilaton spacetime remain largely unexplored. Similarly, while it has been shown that the coherence of the GHZ state is weaker than that of the W state in this context, the comparative nonseparability of these states within the specific framework of the GHS dilaton black hole has not been systematically investigated. Addressing these two critical questions provides another key motivation for our research. Building on these motivations, we will study the impact of the Hawking effect on the nonseparability of N-partite GHZ and W states by using the AR $q$-conditional entropy for both free bosonic and fermionic fields in the background of the GHS dilaton black hole. Initially, we assume that $N$ observers are distributed in an asymptotically flat region, each sharing GHZ and W states. Next, we consider that $N-1$ observers remain stationary in the flat region, while a single observer is positioned near the event horizon of the black hole.
@@ II Abe-Rajagopal $q$-conditional entropy
The computable form of this conditional entropy for any $A-B$ bipartition in a multipartite state can be obtained by substituting Eq.(4) into Eq.(5)
@@
where $\lambda_{n}$ and $\lambda_{m}$ represent the eigenvalues of the whole system $\rho_{AB}$ and one of its subsystem $\rho_{A}$, respectively. This conditional entropy adheres to the nonadditivity relation:$S_{q}(A,B)=S_{q}(A)+S_{q}(B)+(1-q)S_{q}(A)S_{q}(B)$. The AR $q$-conditional entropy yields negative values for nonseparable states, making the positivity of the AR $q$-conditional entropy a criterion for capturing separability in bipartite states. From this point forward, we quantify nonseparability using the AR $q$-conditional entropy for GHZ and W states, with one of the qubits positioned near the event horizon of the black hole.
@@ III Quantization of bosonic and fermionic fields in GHS dilaton spacetime
where $D$ and $M$ are parameters associated with the dilaton field and the mass of the black hole, respectively. The relationship among $D$, the magnetic charge $Q$, and $M$ is described by the equation $D=Q^{2}/2M^{2}$ rfv01 ; rfv02 . In this paper, we adopt natural units where $G=c=\hbar=\kappa_{B}=1$.
@@ III.1 Bosonic field
where $\mu/\nu$ represents the coordinate components in spacetime, $\psi$ is the scalar filed, and $R$ is the Ricci scalar curvature. The coupling between the gravitational field and the scalar field is represented by the term $\xi R\psi$, where $\xi$ is a numerical couple factor. The normal mode solution can be expressed as
@@
where $f(r)=(r-2M)/(r-2D)$ and the tortoise coordinate $r_{\ast}$ is defined as $dr_{\ast}=dr/f(r)$ A5 .
@@
By utilizing the relation $-1=e^{i\pi}$ and ensuring that Eq.(16) is analytic in the lower half-plane of $U$, a complete basis for positive-energy $U$ modes can be expressed as
@@ IV N-partite nonseparability of bosonic and fermionic fields in dilaton spacetime
where the mode $i(i=1,2,\ldots,N)$ is detected by observer $O_{i}$, at the same point in the asymptotically flat region of the dilaton black hole. Here, we use $B$ and $F$ to denote bosonic and fermionic modes, respectively. After sharing their qubits, one observer hovers near the event horizon of the dilaton black hole, while the remaining $N-1$ observers remain stationary in the asymptotically flat region.
@@ IV.1 N-partite nonseparability of bosonic field
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{B,\mathrm{GHZ}}$, corresponding to the modes of the last qubit in exterior region of the dilaton black hole, is obtained by tracing out subsystems of the front $N-1$ particles, i.e., $\rho_{N_{\mathrm{out}}}^{B,\mathrm{GHZ}}=\mathrm{Tr}_{123\ldots N-1}[\rho_{123\ldots N_{\mathrm{out}}}^{B,\mathrm{GHZ}}]$, and is given by
@@
which is a block matrix in the subspace of $\{|10\ldots 0n\rangle,|01\ldots 0n\rangle,\cdots|00\ldots 1n\rangle,|00\ldots 0n+1\rangle\}$. The matrix $\Delta^{\prime}_{n}(\rho_{123\ldots N_{\mathrm{out}}}^{B,\mathrm{W}})$ has a specific block structure. It is a $N\times N$ matrix, where the first $N$-1 rows consist of the same repeated elements $(1-\alpha)\alpha^{n}$ besides the last column of these rows which contains $(1-\alpha)^{\frac{3}{2}}\sqrt{n+1}\alpha^{n}$, and the last row consists entirely of $(1-\alpha)^{\frac{3}{2}}\sqrt{n+1}\alpha^{n}$ for the first N-1 columns with the bottom-right element being $(1-\alpha)^{2}(n+1)\alpha^{n}$. The eigenvalues of this $n$-th block of the state $\rho_{123\ldots N_{out}}^{B,\mathrm{W}}$ are 0 and
@@
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{B,\mathrm{W}}$ is obtained by tracing out $N-1$ subsystems as
@@
$+$ $N|m\rangle_{N,\mathrm{out}}\langle m|\bigg{\}},$
@@
In Fig.1, we plot AR $q$-conditional entropy $S^{\mathrm{GHZ},_{3}}_{B,_{q}}(ABC)$ and $S^{\mathrm{W},_{3}}_{B,_{q}}(ABC)$ as a function of the dilaton $D$ for different $q$, derived from Eqs.(39) and (47). From Fig.1, we observe that $q$-conditional entropy remains negative and increases monotonically with the increase of the dilaton $D$, where more negative values indicate stronger nonseparability. This implies that the nonseparability of both GHZ and W states decreases as the dilaton $D$ increases. Notably, as the dilaton $D$ approaches the mass $M$ of the black hole, $q$-conditional entropy $S^{\mathrm{GHZ},_{3}}_{B,_{q}}(ABC)$ and $S^{\mathrm{W},_{3}}_{B,_{q}}(ABC)$ tend toward zero for all values of any $q$, signaling a transition from nonseparability to separability in the extreme limit $D\rightarrow M$. This marks a significant shift in the entanglement properties of quantum states at this threshold. Furthermore, the $q$-conditional entropy of W state is consistently larger than that of GHZ state within dilaton spacetime, indicating weaker nonseparability for W state under the influence of the Hawking effect. In contrast, quantum coherence of W state surpasses that of GHZ state in the dilaton black hole A16 . This contrast underscores the distinct behaviors of GHZ and W states in relativistic quantum information tasks. When selecting quantum resources for such tasks, the choice between GHZ and W states must be carefully considered, as it significantly impacts the efficiency and outcomes of quantum information protocols in the presence of the dilaton black hole.
@@ IV.2 N-partite nonseparability of fermionic field
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{F,\mathrm{GHZ}}$ is obtained by tracing out subsystems of the front $N-1$ particles, and is given by
@@
$+\sum_{i=1}^{N-1}|0\rangle_{i}\langle 1|\bigotimes_{j=1(j\neq i)}^{N-1}|0\rangle_{j}\langle 0|[e^{-8\pi\omega(M-D)}+1]^{-1}|1\rangle_{N,\mathrm{out}}\langle 0|.$
@@
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{F,\mathrm{W}}$ is obtained by tracing out $N-1$ subsystems as
@@
In Fig.2, we present the $q$-conditional entropy $S^{\mathrm{GHZ},_{3}}_{F,_{q}}(ABC)$ and $S^{\mathrm{W},_{3}}_{F,_{q}}(ABC)$ of fermionic field as a function of the dilaton $D$ for different $q$, calculated from Eqs.(53) and (59). From Fig.2, we can see that the behavior of $S^{\mathrm{GHZ},_{3}}_{F,_{q}}(ABC)$ in fermionic field is similar to that of $S^{\mathrm{GHZ},_{3}}_{B,_{q}}(ABC)$ of bosonic field in curved spacetime. However, a striking distinction emerges with the $q$-conditional entropy $S^{\mathrm{W},_{3}}_{F,_{q}}(ABC)$ of fermionic field, which is a non-monotonic function of the $D$. This non-monotonicity indicates that the Hawking effect from the black hole significantly enhances the nonseparability of W state in fermionic field. As the $D$ increases, the initial nonseparability of W state reaches a maximum, demonstrating that the Hawking effect induces a net increase in nonseparability. Therefore, the properties of $S^{\mathrm{W},_{3}}_{F,_{q}}(ABC)$ of fermionic field are different from those of bosonic field in the dilaton black hole. Unlike the $q$-conditional entropy of bosonic field, as the $D$ approaches the mass $M$ of the black hole, the $q$-conditional entropy of fermionic field tends toward non-zero negative values for all values of $q$, indicating that both GHZ and W states retain their nonseparability in the extreme black hole limit. In other words, at the threshold $D\rightarrow M$, the quantum states of fermionic field still remain entangled, highlighting a fundamental difference between fermionic and bosonic fields in the context of the dilaton black hole.
Furthermore, from both Fig.1 and Fig.2, we observe that the $q$-conditional entropy of bosonic field is consistently greater than that of fermionic field within dilaton spacetime, suggesting that the nonseparability of bosonic field is weaker than that of fermionic field under the influence of the Hawking effect. In contrast, quantum coherence is stronger for bosonic field than for fermionic field in the dilaton black hole A11 ; A16 , emphasizing the distinct quantum properties of these fields in a relativistic background. This highlights the importance of selecting the appropriate quantum resources when conducting relativistic quantum information tasks, as the choice of particle type can significantly influence the efficiency and outcome of quantum protocols.
@@
Through the analytical expressions in Eqs.(39) and (53), we find that the $q$-conditional entropy of GHZ state remains independent of the number of qubits $N$ in both bosonic and fermionic fields. This stands in stark contrast to the behavior observed in W state, where, as seen in Eqs.(47) and (59), the $q$-conditional entropy is explicitly dependent on $N$ in both types of fields. To illustrate this relationship, we plot the AR $q$-conditional entropy $S^{\mathrm{W},_{N}}_{B,_{q}}(N)$ and $S^{\mathrm{W},_{N}}_{F,_{q}}(N)$ as a function of the dilaton $D$ for different $q$ and $N$, as shown in Fig.3. From Fig.3, it is evident that the $q$-conditional entropy of W state increases with the growth of the $N$, indicating a decrease in nonseparability as the number of qubits grows within the dilaton black hole. This trend is indicative of the complex interplay between quantum correlations that emerge as more qubits are introduced into the system, directly influencing the conditional entropy. This behavior is akin to the properties of quantum entanglement in W state under relativistic environments, where quantum entanglement diminishes as the $N$ grows. Interestingly, while the nonseparability of W state weakens with increasing $N$, its quantum coherence exhibits a monotonically increasing trend within the dilaton spacetime A16 . This duality between decreasing entanglement and increasing coherence underscores the nuanced and complex nature of quantum resources in relativistic regimes. These findings emphasize the importance of carefully selecting multipartite quantum resources for relativistic quantum information tasks, as they reveal competing factors that affect the overall quantum information dynamics in curved spacetime.

### v2

@@ I Introduction
From Einstein’s perspective, the gravitational collapse of sufficiently massive stars leads to the formation of black holes, fascinating and profound entities that challenge and enrich our understanding of the universe. In recent years, both indirect and direct astronomical observations have provided strong evidence confirming the existence of these mysterious cosmic objects Q3 ; Q4 ; Q5 ; Q6 ; Q7 ; Q8 ; Q9 ; Q10 ; Q11 . A prominent example of a black hole within this framework is the Garfinkle-Horowitz-Strominger (GHS) dilaton black hole, a solution that arises from the intriguing predictions of string theory rfv01 ; rfv02 ; rfv1 ; rfv2 ; rfv3 . The presence of a dilaton field can significantly alter the properties of black hole geometries, making the study of dilaton black holes crucial for both gravitational theory and quantum mechanics. Due to their immense distances from us and their extraordinary properties, black holes have long remained enveloped in mystery, becoming a central focus of cutting-edge research. One of the most intriguing challenges in this field is the information loss paradox of the black hole Q12 ; rfv4 ; rfv5 ; rfv6 . As we all know, Hawking radiation is believed to be the root cause of the information loss paradox, as it suggests that information falling into a black hole may be irretrievably lost, seemingly violating the principles of quantum mechanics. Quantum properties, particularly in the context of Hawking radiation, may offer a crucial pathway toward resolving this paradox. Moreover, in realistic environments, quantum systems are inevitably influenced by gravitational effects, making it essential to consider the impact of Hawking radiation on quantum resources.
@@
Quantum information in gravitational background R1 ; R2 ; R3 ; R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; BJ1 ; BJ2 ; BJ3 ; R10 ; R11 ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R18 ; RM1 ; RM2 ; RM3 ; RM4 ; RM5 ; RM6 ; RM7 ; RM8 ; RM9 ; RM10 ; RM11 ; RM12 ; QQRM12 ; WDD1 ; WDD2 ; WDD3 ; BJ4 ; WDD4 ; WDD5 ; B31 ; QQQQR8 ; qert1 ; RTYR4QQZ ; qert2 ; qert3 ; qert4 ; qert5 ; qert6 ; qert7 , an interdisciplinary field that bridges quantum information theory, quantum field theory, and general relativity. Theoretically, the influence of the gravitational effect on bipartite steering, entanglement, discord, coherence, quantum Fisher information, and entropic uncertainty relations has been investigated extensively in the background of a dilaton black hole A1 ; A2 ; A3 ; A4 ; A5 ; A6 ; A7 ; A8 ; A9 ; A10 ; A11 ; A12 ; A13 ; A14 ; A15 . As the complexity of quantum information tasks continues to increase, bipartite and tripartite inseparable states can no longer meet the growing demands, necessitating the use of N-partite inseparable states to handle these advanced tasks. For example, Deng et al. have reported new Gaussian boson sampling experiments with pseudo-photon-number-resolving detection, which register up to 255 photon-click events Q13 , and Google quantum AI and collaborators have demonstrated surface code memory below threshold in new Willow architecture by using 105 nonseparable qubits Q14 . It is necessary to understand the N-partite nonseparable quantum systems in gravitational background, because the influence of gravitational effects on realistic quantum systems is a non-negligible factor. However, measuring N-partite entanglement and obtaining its analytical expression in the dilaton black hole, particularly for complex W state, is a challenging task. Thus, one of our primary motivations is to use the Abe-Rajagopal (AR) $q$-conditional entropy to quantify the nonseparability of N-partite systems under the dilaton black hole background. Notably, this approach places stricter constraints on separability compared to traditional methods, such as the von Neumann conditional entropy, offering more nuanced insights into quantum correlations under the extreme gravitational background A17 ; A18 ; A19 .
@@
Previous studies have indicated that fermionic coherence is generally weaker than bosonic coherence in the dilaton black hole background A11 ; A16 . However, the precise differences between the nonseparability of bosonic and fermionic states in dilaton spacetime remain largely unexplored. Similarly, while it has been shown that the coherence of the GHZ state is weaker than that of the W state in this context, the comparative nonseparability of these states within the specific framework of the GHS dilaton black hole has not been systematically investigated. Addressing these two critical questions provides another key motivation for our research. Building on these motivations, we will study the impact of the Hawking effect on the nonseparability of N-partite GHZ and W states by using the AR $q$-conditional entropy for both free bosonic and fermionic fields in the background of the GHS dilaton black hole. Initially, we assume that $N$ observers are distributed in an asymptotically flat region, each sharing GHZ and W states. Next, we consider that N1N-1 observers remain stationary in the flat region, while a single observer is positioned near the event horizon of the black hole.
@@ II Abe-Rajagopal $q$-conditional entropy
The computable form of this conditional entropy for any ABA-B bipartition in a multipartite state can be obtained by substituting Eq.(4) into Eq.(5)
@@
[…]
@@ III Quantization of bosonic and fermionic fields in GHS dilaton spacetime
where $D$ and $M$ are parameters associated with the dilaton field and the mass of the black hole, respectively. The relationship among $D$, the magnetic charge $Q$, and $M$ is described by the equation $D=Q^{2}/2M$ rfv01 ; rfv02 . In this paper, we adopt natural units where $G=c=\hbar=\kappa_{B}=1$.
@@ III.1 Bosonic field
where $\mu/\nu$ represents the coordinate components in spacetime, $\psi$ is the scalar field, and $R$ is the Ricci scalar curvature. The coupling between the gravitational field and the scalar field is represented by the term $\xi R\psi$, where $\xi$ is a numerical couple factor. The normal mode solution can be expressed as
@@
where f(r)=(r2M)/(r2D)f(r)=(r-2M)/(r-2D) and the tortoise coordinate $r_{\ast}$ is defined as $dr_{\ast}=dr/f(r)$ A5 .
@@
By utilizing the relation $1=e^{i\pi}$ and ensuring that Eq.(16) is analytic in the lower half-plane of $U$, a complete basis for positive-energy $U$ modes can be expressed as
@@ IV N-partite nonseparability of bosonic and fermionic fields in dilaton spacetime
where the mode $i(i=1,2,\ldots,N)$ is detected by observer $O_{i}$, at the same point in the asymptotically flat region of the dilaton black hole. Here, we use $B$ and $F$ to denote bosonic and fermionic modes, respectively. After sharing their qubits, one observer hovers near the event horizon of the dilaton black hole, while the remaining N1N-1 observers remain stationary in the asymptotically flat region.
@@ IV.1 N-partite nonseparability of bosonic field
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{B,\mathrm{GHZ}}$, corresponding to the modes of the last qubit in exterior region of the dilaton black hole, is obtained by tracing out subsystems of the front N1N-1 particles, i.e., $\rho_{N_{\mathrm{out}}}^{B,\mathrm{GHZ}}=\mathrm{Tr}_{123\ldots N-1}[\rho_{123\ldots N_{\mathrm{out}}}^{B,\mathrm{GHZ}}]$, and is given by
@@
which is a block matrix in the subspace of $\{|10\ldots 0n\rangle,|01\ldots 0n\rangle,\cdots|00\ldots 1n\rangle,|00\ldots 0n+1\rangle\}$. The matrix $\Delta^{\prime}_{n}(\rho_{123\ldots N_{\mathrm{out}}}^{B,\mathrm{W}})$ has a specific block structure. It is a $N\times N$ matrix, where the upper-left $(N-1)\times(N-1)$ submatrix consists entirely of the same element, $(1-\alpha)\alpha^{n}$. The entries in both the last row and the last column are all equal to $(1-\alpha)^{\frac{3}{2}}\sqrt{n+1}\alpha^{n}$, with the exception of the bottom-right corner element, which is $(1-\alpha)^{2}(n+1)\alpha^{n}$.
The eigenvalues of this $n$-th block of the state $\rho_{123\ldots N_{out}}^{B,\mathrm{W}}$ are 0 and
@@
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{B,\mathrm{W}}$ is obtained by tracing out N1N-1 subsystems as
@@
$+$ $(N-1)|m\rangle_{N,\mathrm{out}}\langle m|\bigg{\}}=\frac{1}{N}e^{-8m\pi\omega(M-D)}[1-e^{-8\pi\omega(M-D)}]\bigg{\{}N-1+me^{8\pi\omega(M-D)}$
$\times$ $[1-e^{-8\pi\omega(M-D)}]\bigg{\}}|m\rangle_{N,\mathrm{out}}\langle m|,$
@@
In Fig.1, we plot AR $q$-conditional entropy $S^{\mathrm{GHZ},_{3}}_{B,_{q}}(ABC)$ and $S^{\mathrm{W},_{3}}_{B,_{q}}(ABC)$ as a function of the dilaton $D$ for different $q$, derived from Eqs.(39) and (47). From Fig.1, we observe that $q$-conditional entropy remains negative and increases monotonically with the increase of the dilaton $D$, where more negative values indicate stronger nonseparability. This implies that the nonseparability of both GHZ and W states decreases as the dilaton $D$ increases. Notably, as the dilaton $D$ approaches the mass $M$ of the black hole, $q$-conditional entropy $S^{\mathrm{GHZ},_{3}}_{B,_{q}}(ABC)$ and $S^{\mathrm{W},_{3}}_{B,_{q}}(ABC)$ tend toward zero for all values of any $q$, signaling a transition from nonseparability to separability in the extreme limit $D\rightarrow M$. This marks a significant shift in the entanglement properties of quantum states at this threshold. Furthermore, the $q$-conditional entropy of W state is consistently larger than that of GHZ state within dilaton spacetime, indicating weaker nonseparability for W state under the influence of the Hawking effect. In contrast, quantum coherence of W state surpasses that of GHZ state in the dilaton black hole A16 . This contrast underscores the distinct behaviors of GHZ and W states in relativistic quantum information tasks. In strong gravitational backgrounds (such as black hole), the Hawking effect can induce decoherence, making the stability of quantum systems crucial during task execution. Although GHZ state provides stronger multipartite entanglement and is suitable for tasks that require high entanglement (such as quantum teleportation and quantum key distribution), their comparatively weaker coherence makes it more susceptible to decoherence in strong gravitational fields. In contrast, W state, despite possessing weaker entanglement, exhibits stronger coherence and is therefore more resistant to gravity-induced decoherence. As a result, in practical relativistic quantum information tasks, it is essential to choose appropriate quantum resources based on specific task requirements: GHZ state is preferable for high-entanglement tasks, while W state is better suited for scenarios demanding stronger noise resistance. This finding highlights the importance of carefully balancing and selecting quantum resources in curved spacetime to enhance the performance and robustness of quantum protocols.
@@ IV.2 N-partite nonseparability of fermionic field
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{F,\mathrm{GHZ}}$ is obtained by tracing out subsystems of the front N1N-1 particles, and is given by
@@
$+\sum_{i=1}^{N-1}|0\rangle_{i}\langle 1|\bigotimes_{j=1(j\neq i)}^{N-1}|0\rangle_{j}\langle 0|[e^{-8\pi\omega(M-D)}+1]^{-\frac{1}{2}}|1\rangle_{N,\mathrm{out}}\langle 0|.$
@@
The reduced density matrix $\rho_{N_{\mathrm{out}}}^{F,\mathrm{W}}$ is obtained by tracing out N1N-1 subsystems as
@@
In Fig.2, we present the $q$-conditional entropy $S^{\mathrm{GHZ},_{3}}_{F,_{q}}(ABC)$ and $S^{\mathrm{W},_{3}}_{F,_{q}}(ABC)$ of fermionic field as a function of the dilaton $D$ for different $q$, calculated from Eqs.(53) and (59). From Fig.2, we can see that the behavior of $S^{\mathrm{GHZ},_{3}}_{F,_{q}}(ABC)$ in fermionic field is similar to that of $S^{\mathrm{GHZ},_{3}}_{B,_{q}}(ABC)$ of bosonic field in curved spacetime. However, a striking distinction emerges with the $q$-conditional entropy $S^{\mathrm{W},_{3}}_{F,_{q}}(ABC)$ of fermionic field, which is a non-monotonic function of the $D$. This non-monotonicity indicates that the Hawking effect from the black hole significantly enhances the nonseparability of W state in fermionic field. As the $D$ increases, the initial nonseparability of W state reaches a maximum, demonstrating that the Hawking effect induces a net increase in nonseparability. Therefore, the properties of $S^{\mathrm{W},_{3}}_{F,_{q}}(ABC)$ of fermionic field are different from those of bosonic field in the dilaton black hole. Unlike the $q$-conditional entropy of bosonic field, as the $D$ approaches the mass $M$ of the black hole, the $q$-conditional entropy of fermionic field tends toward non-zero negative values for all values of $q$, indicating that both GHZ and W states retain their nonseparability in the extreme black hole limit. In other words, at the threshold $D\rightarrow M$, the quantum states of fermionic field still remain entangled, highlighting a fundamental difference between fermionic and bosonic fields in the context of the dilaton black hole. Furthermore, from both Fig.1 and Fig.2, we observe that the $q$-conditional entropy of bosonic field is consistently greater than that of fermionic field within dilaton spacetime, suggesting that the nonseparability of bosonic field is weaker than that of fermionic field under the influence of the Hawking effect. In contrast, quantum coherence is stronger for bosonic field than for fermionic field in the dilaton black hole A11 ; A16 , emphasizing the distinct quantum properties of these fields in a relativistic background. Due to the differences between the Bose-Einstein and Fermi-Dirac distributions, bosons and fermions exhibit fundamentally distinct behaviors in the background of a dilaton black hole. For instance, in the extreme black hole limit, entanglement for bosons vanishes completely, whereas fermionic entanglement can partially survive, demonstrating greater resilience. As a result, fermionic nonseparability is generally more robust than that of bosons in such spacetime. The quantum coherence of multiple particles reflects the ability of each particle to maintain superposition within its local state space, while quantum entanglement refers to the non-classical correlations between particles, such that the total state cannot be factorized into a product of individual subsystems. Owing to their fundamentally different physical characteristics, it is found that fermionic coherence is consistently smaller than bosonic coherence in the dilaton black hole. This highlights the importance of selecting the appropriate quantum resources when conducting relativistic quantum information tasks, as the choice of particle type can significantly influence the efficiency and outcome of quantum protocols.
@@
Through the analytical expressions in Eqs.(39) and (53), we find that the $q$-conditional entropy of GHZ state remains independent of the number of qubits $N$ in both bosonic and fermionic fields. This stands in stark contrast to the behavior observed in W state, where, as seen in Eqs.(47) and (59), the $q$-conditional entropy is explicitly dependent on $N$ in both types of fields. To illustrate this relationship, we plot the AR $q$-conditional entropy $S^{\mathrm{W},_{N}}_{B,_{q}}(N)$ and $S^{\mathrm{W},_{N}}_{F,_{q}}(N)$ as a function of the dilaton $D$ for different $q$ and $N$, as shown in Fig.3. From Fig.3, it is evident that the $q$-conditional entropy of W state increases with the growth of the $N$, indicating a decrease in nonseparability as the number of qubits grows within the dilaton black hole. This trend is indicative of the complex interplay between quantum correlations that emerge as more qubits are introduced into the system, directly influencing the conditional entropy. This behavior is akin to the properties of quantum entanglement in W state under relativistic environments, where quantum entanglement diminishes as the $N$ grows. Interestingly, while the nonseparability of W state weakens with increasing $N$, its quantum coherence exhibits a monotonically increasing trend within the dilaton spacetime A16 . This is because, as the number of particles $N$ in the W state increases, a fixed amount of quantum correlations is distributed over more subsystems, leading to a weakening of the overall entanglement (nonseparability), especially in pairwise correlations. However, since the W state involves superpositions of more basis states, its quantum coherence increases. In the background of the GHS dilaton black hole, the thermal noise induced by the Hawking effect more severely degrades entanglement, while coherence exhibits greater robustness against such noise-it can not only survive but may even increase with growing $N$. This duality between decreasing entanglement and increasing coherence underscores the nuanced and complex nature of quantum resources in relativistic regimes. These findings emphasize the importance of carefully selecting multipartite quantum resources for relativistic quantum information tasks, as they reveal competing factors that affect the overall quantum information dynamics in curved spacetime.